# Magic dynamics in many-body localized systems

### v2

@@ II Details on the $\ell$-bit model
In the MBL regime, when the disorder is coupled to on-site operators 111 If the disorder is not coupled to on-site operators, the standard $\ell$-bit description may fail to capture the system’s physics. In such cases, a more sophisticated RSRG-X approach is required to correctly describe the model’s properties [Aramthotill24], the system is simply described by the emergence of an extensive large set of integrals of motion, also known as localized bits ($\ell$-bits). The quasilocal structure of the $\ell$-bits encapsulates key signatures of MBL, making the Hamiltonian $\hat{\mathcal{H}}_{\mathrm{\ell-bit}}$ (see Eq. 3 of the main letter) a natural framework to probe the nonstabilizerness dynamics in MBL. The physical degrees of freedom are related to the $\ell$-bits by quasilocal unitary transformations $U$, satisfying $\hat{\tau}_{j}^{z}=U\hat{Z}_{j}U^{\dagger}$. Ideally, $U$ reflects the system’s localized nature, with its matrix elements decaying exponentially with the distance between $\ell$-bits. However, for practical purposes, a good approximation is to take $U=I$, where $I$ is the identity operator, significantly simplifying the analysis while retaining the essential features of MBL dynamics.
@@
In Fig. S.3(a), we present the time evolution of the entanglement entropy for different localization strengths $\alpha$. Although $\mathcal{S}$ exhibits the expected logarithmic growth over time, its saturation value remains well below the Page value (black dotted line), indicating that the spins are not fully dephased. The extent of this saturation depends on $\alpha$. A similar dependence on $\alpha$ is observed in the saturation value of $\mathcal{M}_{2}$. However, as shown in Fig. S.3(b), the overall time-scaling behavior of $\mathcal{M}_{2}$ remains unchanged, suggesting that while localization strength influences the final magic content, the fundamental growth dynamics are robust to variations in $\alpha$.
III Entanglement as the “internal-clock” for the nonstabilizerness dynamics

### v3

@@ I Details on the $\ell$-bit model
In the MBL regime, when the disorder is coupled to on-site operators  [^1] If the disorder is not coupled to on-site operators, the standard $\ell$-bit description may fail to capture the system’s physics. In such cases, a more sophisticated RSRG-X approach is required to correctly describe the model’s properties [141], the system is simply described by the emergence of an extensive large set of integrals of motion, also known as localized bits ($\ell$-bits). The quasilocal structure of the $\ell$-bits encapsulates key signatures of MBL, making the Hamiltonian $\hat{\mathcal{H}}_{\mathrm{\ell-bit}}$ (see Eq. 3 of the main letter) a natural framework to probe the nonstabilizerness dynamics in MBL. The physical degrees of freedom are related to the $\ell$-bits by quasilocal unitary transformations $U$, satisfying $\hat{\tau}_{j}^{z}=U\hat{Z}_{j}U^{\dagger}$. Ideally, $U$ reflects the system’s localized nature, with its matrix elements decaying exponentially with the distance between $\ell$-bits. However, for practical purposes, a good approximation is to take $U=I$, where $I$ is the identity operator, significantly simplifying the analysis while retaining the essential features of MBL dynamics.
@@
In Fig. S.3(a), we present the time evolution of the entanglement entropy for different localization strengths $\alpha$. Although $\mathcal{S}$ exhibits the expected logarithmic growth over time, its saturation value remains well below the Page value (black dotted line), indicating that the spins are not fully dephased. The extent of this saturation depends on $\alpha$. A similar dependence on $\alpha$ is observed in the saturation value of $\mathcal{M}_{2}$. However, as shown in Fig. S.3(b), the overall time-scaling behavior of $\mathcal{M}_{2}$ remains unchanged, suggesting that while localization strength influences the final nonstabilizerness content, the fundamental growth dynamics are robust to variations in $\alpha$.